OpenStudy (aaronandyson):

Find the interior angle for an N sided polygon. Also find the sum of all interior angle of the polygon.

OpenStudy (aaronandyson):

@hartnn

OpenStudy (aaronandyson):

But how do I apply it?

OpenStudy (aaronandyson):

@mayankdevnani

OpenStudy (mayankdevnani):

acha chal octagon ke liye dekhte h,okay ?

OpenStudy (aaronandyson):

muhje bas ye q karna h "polygon" is not enuf info octa hi lelete hain

OpenStudy (mayankdevnani):

exterior angle nikalne ka simple tarika h \[\large \bf \frac{180}{n}\]

OpenStudy (mayankdevnani):

where, n=number of sides of polygon

OpenStudy (mayankdevnani):

theek h ?

OpenStudy (aaronandyson):

haan

OpenStudy (aaronandyson):

@mayankdevnani

OpenStudy (mayankdevnani):

interior angle ke liye :- \[\large \bf 180-n\]

OpenStudy (aaronandyson):

too sum kaise nikale?

OpenStudy (mayankdevnani):

formula hota h \[\large \bf 180(n-2)\]

OpenStudy (aaronandyson):

hume "n" hi ni malum ;-;

OpenStudy (aaronandyson):

@Ashleyisakitty @johnweldon1993 @sleepyjess @Kainui

OpenStudy (baru):

here is an interesting property of any polygon. The sum of its exterior angles is 360 always. so if a regular polygon has 'n' sides, each exterior angles measures 360/n ok?

OpenStudy (aaronandyson):

Got that!

OpenStudy (baru):

now, interior angle + exterior angle = 180 so we have interior angle = 180 - exterior angle or interior angle=180- (360/n) ok?

OpenStudy (aaronandyson):

Okay.

OpenStudy (baru):

this is the answer for each interior angle interior angle=180- (360/n) a polygon with 'n' sides has 'n' interior angles, so sum of interior angles = interior angle \(\times\) n =180n-360

OpenStudy (aaronandyson):

n cancels out?

OpenStudy (aaronandyson):

No,wait. You took the LCM and then multiplied it with n?

OpenStudy (aaronandyson):

\[\frac{ 180n*360 }{ n }*n\]

OpenStudy (baru):

yep

OpenStudy (aaronandyson):

180n - 360??

OpenStudy (baru):

yes, 180n-360 is the sum of interior angles

OpenStudy (aaronandyson):

and each interior angle is 180n-360 all over n?

OpenStudy (baru):

yep

OpenStudy (aaronandyson):

That's it?

OpenStudy (baru):

thats it

OpenStudy (aaronandyson):

Can you help me with a few more question? And thanks.

OpenStudy (baru):

yea sure